The average age of a group of teachers and students is 20. The average age of the teachers is 35. The average age of the students is 15. I need to know the ratio of teachers to students and have it answered as a fraction in simplest form.
Answer by scott8148(6628)

(35t + 15s) / (t + s) = 20

35t + 15s = 20t + 20s

15t = 5s

t/s = 5/15 = 1/3